# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ 2.1. Basic Framework
(10) $\mathcal{E}^{*}_{\mathbb{r}}:=\underset{\mathbb{X}\in\mathcal{V}_{\mathbb{r}}}{\min}\;\mathcal{j}(\mathbb{X}).$
Since the tensor contract mapping $\tau\colon\mathcal{U}_{\mathbb{r}}\rightarrow\mathcal{M}_{\mathbb{r}}$ is a submersion, it readily follows that two minimization problems (6) and (10) are equivalent in the sense that $E^{*}_{\mathbb{r}}=\mathcal{E}^{*}_{\mathbb{r}}$, and if $\mathbb{X}$ is a local minimizer of the energy functional $\mathcal{j}$ in $\mathcal{V}_{\mathbb{r}}$, then $\tau(\mathbb{X})$ is a local minimizer of $\mathcal{J}$ in $\mathcal{T}_{\mathbb{r}}$ and vice-versa. It is important to note, however, that since the tensor contraction mapping $\tau$ is not injective, even if the minimizer $\tau(\mathbb{X})$ of $\mathcal{J}$ is locally unique, the corresponding minimizer $\mathbb{X}$ of $\mathcal{j}$ is not.
@@ 2.2. Alternating Minimization on the Tensor Train Parameter Space
The goal of this section is to briefly describe the celebrated alternating optimization algorithms used to numerically resolve the minimization problem (10) on the tensor train parameter space $\overline{\mathcal{U}}_{\mathbb{r}}$. For energy functionals of the form (2), these algorithms are known as the one-site and two-site DMRG methods. Throughout this section, we will make use of the notions introduced in Section 2.1.
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$ and $r_{0}=r_{d}=1$, let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, let $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined according to Definition 4, let $\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$, let $j\in\{1,\ldots,d\}$, and let $k\in\{1,\ldots,d-1\}$. Then we define the one-site retraction operator $\mathbb{P}_{\mathbb{X},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ as the mapping such that for all $\mathbb{V}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ it holds that
$\mathbb{P}_{\mathbb{X},j,1}\mathbb{V}_{j}=\tau\big{(}\mathbb{X}_{1},\ldots,\mathbb{X}_{j-1},\mathbb{V}_{j},\mathbb{X}_{j+1},\ldots,\mathbb{X}_{d}\big{)}.$
Additionally, we define the two-site retraction operator $\mathbb{P}_{\mathbb{X},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ as the mapping such that for all $\mathbb{W}_{k,k+1}\in\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}$, all $i\in\{1,\ldots,d\}$ and all ${x}_{i}\in\{1,\ldots,n_{i}\}$ it holds that
$\big{(}\mathbb{P}_{\mathbb{X},k,2}\mathbb{W}_{k,k+1}\big{)}(\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{d})$
$=$ $\mathbb{X}_{1}(:,x_{1},:)\ldots\mathbb{X}_{k-1}(:,x_{k-1},:)\mathbb{W}_{k,k+1}(:,x_{k},x_{k+1},:)\mathbb{X}_{k+2}(:,x_{k+2},:)\ldots\mathbb{X}_{d}(:,x_{d},:).$
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$ and $r_{0}=r_{d}=1$, let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, let $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined according to Definition 4, let $j\in\{1,\ldots,d\}$ and $k\in\{1,\ldots,d-1\}$, let $\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$ and $\mathbb{Y}=(\mathbb{Y}_{1},\ldots,\mathbb{Y}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$ be a $j$-orthogonal and $k$-orthogonal tensor train decomposition respectively, and let the one-site and two-site retraction operators $\mathbb{P}_{\mathbb{X},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ and $\mathbb{P}_{\mathbb{Y},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ respectively, be defined according to Definition 7. Then the operators $\mathbb{P}_{\mathbb{X},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ and $\mathbb{P}_{\mathbb{Y},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ are orthogonal projections in the sense that $\forall\mathbb{U}_{j},\mathbb{V}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ and $\forall\mathbb{W}_{k,k+1},\mathbb{Z}_{k,k+1}\in\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}$ it holds that
$\big{\langle}\mathbb{P}_{\mathbb{X},j,1}\mathbb{U}_{j},\mathbb{P}_{\mathbb{X},j,1}\mathbb{U}_{j}\big{\rangle}$ $=\big{\langle}\mathbb{U}_{j},\mathbb{U}_{j}\big{\rangle},\quad\text{and}$
$\big{\langle}\mathbb{P}_{\mathbb{Y},k,2}\mathbb{W}_{k,k+1},\mathbb{P}_{\mathbb{Y},k,2}\mathbb{Z}_{k,k+1}\big{\rangle}$ $=\big{\langle}\mathbb{W}_{k,k+1},\mathbb{Z}_{k,k+1}\big{\rangle}.$
@@
See [18, Lemma 3.2]. ∎
Lemma 8 indicates that the one-site and two-site retraction operators reduce to orthogonal mappings if the underlying tensor train decomposition satisfies the appropriate orthogonality condition. This fact has several useful numerical consequences in the practical implementation of the one-site and two-site DMRG methods, which explains why these algorithms typically involve an orthogonalization step wherein the underlying tensor train decomposition is appropriately orthogonalized prior to the optimization step. To explain this process further, we next introduce the notion of the one-site and two-site DMRG micro-iteration.
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$ and $r_{0}=r_{d}=1$, let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, let $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined according to Definition 4, let $\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big{(}\mathcal{T}_{\mathbb{r}}\cap U\big{)}\subset{\mathcal{U}}_{\mathbb{r}}$, let $j\in\{1,\ldots,d\}$, let $k\in\{1,\ldots,d-1\}$, and for a given $\mathbb{X}\in\mathcal{U}_{\mathbb{r}}$ let the retraction operators $\mathbb{P}_{\mathbb{X},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ and $\mathbb{P}_{\mathbb{X},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ be defined according to Definition 7. Then we define the $j^{\rm th}$ one-site DMRG micro-iteration $\mathcal{S}_{j}\colon\mathcal{V}_{\mathbb{r}}\rightarrow\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ as the mapping with the property that
(11) $\forall\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})\in{\mathcal{V}}_{\mathbb{r}}\colon\qquad\mathcal{S}_{j}(\mathbb{X})=\underset{\mathbb{Z}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}}{\text{\rm argmin}}\mathcal{J}\circ\mathbb{P}_{\mathbb{X},j,1}.$
Additionally, we define the $k^{\rm th}$ two-site DMRG micro-iteration denoted $\mathcal{S}_{k,k+1}\colon\mathcal{V}_{\mathbb{r}}\rightarrow\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}$ as the mapping with the property that
(12) $\forall\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})\in{\mathcal{V}}_{\mathbb{r}}\colon\qquad\mathcal{S}_{k,k+1}(\mathbb{X})=\underset{\mathbb{Z}_{k,k+1}\in\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}}{\text{\rm argmin}}\mathcal{J}\circ\mathbb{P}_{\mathbb{X},k,2}.$
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$ and $r_{0}=r_{d}=1$, let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, let $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined according to Definition 4, let $\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big{(}\mathcal{T}_{\mathbb{r}}\cap U\big{)}\subset{\mathcal{U}}_{\mathbb{r}}$, let $j\in\{1,\ldots,d\}$, let $k\in\{1,\ldots,d-1\}$,let $\mathbb{X}$ and $\mathbb{Y}\in\mathcal{V}_{\mathbb{r}}$ be a $j$-orthogonal and $k$-orthogonal tensor train decomposition respectively, let the one-site and two-site retraction operators $\mathbb{P}_{\mathbb{X},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ and $\mathbb{P}_{\mathbb{X},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ be defined as in Definition 7, and assume that the underlying energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}$ is given by Equation (2), i.e., $\mathcal{J}(\mathbb{U})=\frac{\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle}{\|\mathbb{U}\|^{2}}$ for all $\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}$ and some symmetric operator $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$. Then the one-site and two-site DMRG micro-iterations (see Equations (11), (12)) can be obtained by solving the eigenvalue problems
$\mathbb{P}_{\mathbb{X},j,1}^{*}\mathbb{A}\mathbb{P}_{\mathbb{X},j,1}\mathbb{W}_{j}=\lambda_{j,\mathbb{A}}\mathbb{W}_{j},\qquad\text{and}\qquad\mathbb{P}_{\mathbb{Y},k,2}^{*}\mathbb{A}\mathbb{P}_{\mathbb{Y},k,2}\mathbb{Z}_{k,k+1}=\widetilde{\lambda}_{k,\mathbb{A}}\mathbb{Z}_{k,k+1},$
where $\lambda_{j,\mathbb{A}}$ and $\widetilde{\lambda}_{k,\mathbb{A}}$ are the smallest eigenvalues of the respective operators.
@@
Lemma 10 implies that for Rayleigh quotient-based energy functionals defined according to Equation (2), each DMRG micro-iteration consists of solving an eigenvalue problem. In the case of one-site DMRG for instance, one must find the smallest eigenvalue of the operator $\mathbb{P}_{\mathbb{X},j,1}^{*}\mathbb{A}\mathbb{P}_{\mathbb{X},j,1}$. It is important to note however that this is true only if the underlying tensor train decomposition $\mathbb{X}\in\mathcal{U}_{\mathbb{r}}$ is $j$-orthogonal. Indeed, if $\mathbb{X}$ is not $j$-orthogonal, then the one-site DMRG micro-iteration will consist of solving instead a generalized eigenvalue problem of the form
$\mathbb{P}_{\mathbb{X},j,1}^{*}\mathbb{A}\mathbb{P}_{\mathbb{X},j,1}\mathbb{V}_{j}=\lambda_{j,\mathbb{A}}\mathbb{B}_{j}^{\rm 1site}\mathbb{V}_{j}\qquad\text{in }\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}},$
where $\mathbb{B}_{j}^{\rm 1site}\colon\mathbb{R}^{r_{j-1}n_{j}r_{j}}\rightarrow\mathbb{R}^{r_{j-1}n_{j}r_{j}}$ is a badly conditioned matrix depending on $\mathbb{X}$. This observation justifies our prior contention that practical implementations of DMRG typically involve an orthogonalization step wherein the underlying tensor train decomposition is appropriately orthogonalized prior to the optimization step.
We are now ready to state the one-site and two-site DMRG algorithms.
@@
Right-orthogonal TT decomposition $\mathbb{X}_{\rm LR}^{0,(0)}\in\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big{(}\mathcal{T}_{\mathbb{r}}\cap U\big{)}\subset{\mathcal{U}}_{\mathbb{r}}$.
@@
$\hskip 5.69046pt\text{Compute either }\mathbb{V}_{i}:=\mathcal{S}_{i}(\mathbb{X}_{\rm LR}^{i,(n+1)})\text{ or }\mathbb{W}_{i,i+1}=\mathcal{S}_{i,i+1}(\mathbb{X}_{\rm LR}^{i,(n+1)}).$
@@
$\hskip 41.25641pt\text{Compute }\mathbb{V}_{i}^{<2>}=\mathbb{Q}_{i}\mathbb{R}_{i}~{}\text{with }\mathbb{Q}_{i}\in\mathbb{R}^{r_{i-1}n_{i}\times r_{i}},~{}\mathbb{R}_{i}\in\mathbb{R}^{r_{i}\times r_{i}},$
@@
$\hskip 41.25641pt\mathbb{X}_{\rm LR}^{i,(n+1)}=(\mathbb{X}^{(n+1)}_{1},\ldots,\mathbb{X}^{(n+1)}_{i-1},\widetilde{\mathbb{X}}_{i}^{(n)},\mathbb{X}^{(n)}_{i+1},\ldots,\mathbb{X}^{(n)}_{d})$
$\hskip 24.18501pt\hookrightarrow\mathbb{X}_{\rm LR}^{i+1,(n+1)}=(\mathbb{X}^{(n+1)}_{1},\ldots,\mathbb{X}^{(n+1)}_{i-1},\mathbb{X}^{(n+1)}_{i},\widetilde{\mathbb{X}}^{(n)}_{i+1},\ldots,\mathbb{X}^{(n)}_{d}),$
$\hskip 41.25641pt\text{where }\mathbb{X}^{(n+1)}_{i}=\mathbb{Q}_{i},~{}\widetilde{\mathbb{X}}^{(n+1)}_{i-1}=\mathbb{R}_{i}\mathbb{X}^{(n)}_{i+1}.$
@@
$\hskip 41.25641pt\text{Compute }\mathbb{W}_{i,i+1}^{<2>}=\mathbb{U}_{i}\mathbb{S}_{i}\mathbb{V}^{*}_{i}~{}\text{with }\mathbb{U}_{i}\in\mathbb{R}^{r_{i-1}n_{i}\times\widetilde{r}_{i}},~{}\mathbb{V}_{i}^{*}\in\mathbb{R}^{\widetilde{r}_{i}\times n_{i+1}\times r_{i+1}},$

### v2

@@ 2.1. Basic Framework
(10) $\mathcal{E}^{*}_{\mathbb{r}}:=\underset{\mathbb{U}\in\mathcal{V}_{\mathbb{r}}}{\min}\;\mathcal{j}(\mathbb{U}).$
Since the tensor contract mapping $\tau\colon\mathcal{U}_{\mathbb{r}}\rightarrow\mathcal{T}_{\mathbb{r}}$ is a submersion, it readily follows that two minimization problems (6) and (10) are equivalent in the sense that $\mathcal{E}^{*}_{\mathbb{r}}=E^{*}_{\mathbb{r}}$, and if $\mathbb{U}$ is a local minimizer of the energy functional $\mathcal{j}$ in $\mathcal{V}_{\mathbb{r}}$, then $\tau(\mathbb{U})$ is a local minimizer of $\mathcal{J}$ in $\mathcal{T}_{\mathbb{r}}$ and vice-versa. It is important to note, however, that since the tensor contraction mapping $\tau$ is not injective, even if the minimizer $\tau(\mathbb{U})$ of $\mathcal{J}$ is locally unique, the corresponding minimizer $\mathbb{U}$ of $\mathcal{j}$ is not.
@@ 2.2. Alternating Minimization on the Tensor Train Parameter Space
The goal of this section is to briefly describe the celebrated alternating optimization algorithms used to numerically resolve the minimization problem (10) on the tensor train parameter space $\overline{\mathcal{U}}_{\mathbb{r}}$. For energy functionals of the form (2), these algorithms are known as the one-site and two-site DMRG methods. Throughout this section, we will make use of the notions introduced in Section 2.1. In particular, we assume that $\mathbb{r}\in\mathbb{N}^{d-1}$ with $r_{0}=r_{d}=1$, that $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, and the space $\overline{\mathcal{U}}_{\mathbb{r}}$ and its non-empty subset $\mathcal{U}_{\mathbb{r}}$ are defined according to Definition 4. Moreover, $j\in\{1,\ldots,d\}$ and $k\in\{1,\ldots,d-1\}$ will denote two arbitrary natural numbers.
@@
Let $\mathbb{U}=(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$. We define the one-site retraction operator $\mathbb{P}_{\mathbb{U},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ as the mapping with the property that for all $\mathbb{V}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ it holds that
$\mathbb{P}_{\mathbb{U},j,1}\mathbb{V}_{j}=\tau\big(\mathbb{U}_{1},\ldots,\mathbb{U}_{j-1},\mathbb{V}_{j},\mathbb{U}_{j+1},\ldots,\mathbb{U}_{d}\big).$
Additionally, we define the two-site retraction operator $\mathbb{P}_{\mathbb{U},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ as the mapping such that for all $\mathbb{W}_{k,k+1}\in\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}$, all $i\in\{1,\ldots,d\}$ and all ${x}_{i}\in\{1,\ldots,n_{i}\}$ it holds that
$\big(\mathbb{P}_{\mathbb{U},k,2}\mathbb{W}_{k,k+1}\big)({x}_{1},{x}_{2},\ldots,{x}_{d})$
$=$ $\mathbb{U}_{1}(:,x_{1},:)\ldots\mathbb{U}_{k-1}(:,x_{k-1},:)\mathbb{W}_{k,k+1}(:,x_{k},x_{k+1},:)\mathbb{U}_{k+2}(:,x_{k+2},:)\ldots\mathbb{U}_{d}(:,x_{d},:).$
@@
Let $\mathbb{U}=(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$ and $\mathbb{V}=(\mathbb{V}_{1},\ldots,\mathbb{V}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$ be a $j$-orthogonal and $k$-orthogonal tensor train decomposition respectively. Then the associated one-site retraction operator $\mathbb{P}_{\mathbb{U},j,1}\colon\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ and two-site retraction operator $\mathbb{P}_{\mathbb{V},k,2}\colon\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\rightarrow\mathbb{R}^{n_{1}\times\ldots\times n_{d}}$ are orthogonal projections in the sense that $\forall\mathbb{W}_{j},\widetilde{\mathbb{W}}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ and $\forall\mathbb{W}_{k,k+1},\widetilde{\mathbb{W}}_{k,k+1}\in\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}$ it holds that
$\big\langle\mathbb{P}_{\mathbb{U},j,1}\mathbb{W}_{j},\mathbb{P}_{\mathbb{U},j,1}\widetilde{\mathbb{W}}_{j}\big\rangle$ $=\big\langle\mathbb{W}_{j},\widetilde{\mathbb{W}}_{j}\big\rangle,\quad\text{and}$
$\big\langle\mathbb{P}_{\mathbb{V},k,2}\mathbb{W}_{k,k+1},\mathbb{P}_{\mathbb{V},k,2}\widetilde{\mathbb{W}}_{k,k+1}\big\rangle$ $=\big\langle\mathbb{W}_{k,k+1},\widetilde{\mathbb{W}}_{k,k+1}\big\rangle.$
@@
See [19, Lemma 3.1] and the preceding discussion. ∎
Lemma 8 has several useful numerical consequences in the practical implementation of DMRG methods and explains, in particular, why DMRG algorithms typically involve an orthogonalization step wherein the underlying tensor train decomposition is appropriately orthogonalized prior to the optimization step. To explain this further, we next introduce the notion of the one-site and two-site DMRG micro-iteration.
@@
Let $\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big(\mathcal{T}_{\mathbb{r}}\cap U\big)\subset{\mathcal{U}}_{\mathbb{r}}$. We define the $j^{\rm th}$ one-site DMRG micro-iteration $\mathcal{S}_{j}\colon\mathcal{V}_{\mathbb{r}}\rightarrow\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ as the mapping with the property that
(11) $\forall\mathbb{U}=(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})\in\mathcal{V}_{\mathbb{r}}\colon\qquad\mathcal{S}_{j}(\mathbb{U})=\underset{\mathbb{W}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}}{\text{\rm argmin}}\mathcal{J}\circ\mathbb{P}_{\mathbb{U},j,1}(\mathbb{W}_{j}).$
Additionally, we define the $k^{\rm th}$ two-site DMRG micro-iteration, denoted $\mathcal{S}_{k,k+1}\colon\mathcal{V}_{\mathbb{r}}\rightarrow\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}$, as the mapping with the property that
(12) $\forall\mathbb{V}=(\mathbb{V}_{1},\ldots,\mathbb{V}_{d})\in{\mathcal{V}}_{\mathbb{r}}\colon\hskip 5.69054pt\mathcal{S}_{k,k+1}(\mathbb{V})=\hskip-8.53581pt\underset{\begin{subarray}{c}\mathbb{W}_{k,k+1}\in\mathbb{R}^{r_{k-1}\times n_{k}\times n_{k+1}\times r_{k+1}}\end{subarray}}{\text{\rm argmin}}\mathcal{J}\circ\mathbb{P}_{\mathbb{V},k,2}(\mathbb{W}_{k,k+1}).$
@@
Let $\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big(\mathcal{T}_{\mathbb{r}}\cap U\big)\subset{\mathcal{U}}_{\mathbb{r}}$, let $\mathbb{U}$ and $\mathbb{V}\in\mathcal{V}_{\mathbb{r}}$ be a $j$-orthogonal and $k$-orthogonal tensor train decomposition respectively, and assume that the underlying energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}$ is given by Equation (2), i.e., $\mathcal{J}(\mathbb{X})=\frac{\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle}{\|\mathbb{X}\|^{2}}$ for all $\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}$ and some symmetric operator $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$. Then the one-site and two-site DMRG micro-iterations can be obtained by solving the eigenvalue problems
$\mathbb{P}_{\mathbb{U},j,1}^{*}\mathbb{A}\mathbb{P}_{\mathbb{U},j,1}\mathbb{W}_{j}=\lambda_{j}^{\mathbb{A}}\mathbb{W}_{j},\qquad\text{and}\qquad\mathbb{P}_{\mathbb{V},k,2}^{*}\mathbb{A}\mathbb{P}_{\mathbb{V},k,2}\mathbb{W}_{k,k+1}={\lambda}_{k,k+1}^{\mathbb{A}}\mathbb{W}_{k,k+1},$
where $\mathbb{P}_{\mathbb{U},j,1}^{*}$ and $\mathbb{P}_{\mathbb{V},k,2}^{*}$ are the adjoints of the retraction operators $\mathbb{P}_{\mathbb{U},j,1}$ and $\mathbb{P}_{\mathbb{V},k,2}$ respectively, and $\lambda_{j}^{\mathbb{A}},{\lambda}_{k,k+1}^{\mathbb{A}}$ are the smallest eigenvalues of the respective operators.
@@
Consider the setting of Lemma 10. As discussed in [19], if $\mathbb{U},\mathbb{V}\in\mathcal{V}_{\mathbb{r}}$ are not $j$-orthogonal and $k$-orthogonal respectively, the DMRG micro-iterations will consist of solving badly-conditioned generalized eigenvalue problems– hence the importance of the orthogonalization step in practical DMRG implementations.
The classical one-site and two-site DMRG algorithms are now presented below in the form of Algorithm 1. We end this section with the following important remark on the well-posedness of successive DMRG micro-iterations.
Remark 12 (Well-Posedness of Successive DMRG Micro-iterations).
In order for the DMRG Algorithm 1 to be well-posed, we require that for a given initialization $\mathbb{U}_{\rm LR}^{0,(0)}\in\mathcal{V}_{\mathbb{r}}$, each subsequent iterate $\mathbb{U}_{\rm LR}^{j,(n)}$, $\mathbb{U}_{\rm RL}^{k,(n)}$  $j\in\{1,\ldots,d-1\},~k\in\{2,\ldots,d\}$, and $n\in\mathbb{N}$ should also be an element of $\mathcal{V}_{\mathbb{r}}$. This is indeed true for quadratic and Rayleigh quotient-based energy functionals $\mathcal{J}$ (and in the two-site DMRG case, for a judicious choice of rank truncation parameter $\widetilde{\mathbb{r}}$) provided that the initialization is sufficiently close to the true minimizer [38, 44]. For a convergence analysis of the one-site DMRG algorithm in these cases, we refer interested readers to [38, 44].
@@
Objective: Compute $\underset{\mathbb{U}\in{\mathcal{V}}_{\mathbb{r}}}{\min}\;\mathcal{j}(\mathbb{U})$ for some choice of $\mathbb{r}\in\mathbb{N}^{d-1}$.
Right-orthogonal TT decomposition $\mathbb{U}_{\rm LR}^{0,(0)}\in\mathcal{V}_{\mathbb{r}}=\tau^{-1}\big(\mathcal{T}_{\mathbb{r}}\cap U\big)\subset{\mathcal{U}}_{\mathbb{r}}$.
@@
$\hskip 5.69046pt\text{Compute either }\mathbb{W}_{i}:=\mathcal{S}_{i}(\mathbb{U}_{\rm LR}^{i,(n+1)})\text{ or }\mathbb{W}_{i,i+1}=\mathcal{S}_{i,i+1}(\mathbb{U}_{\rm LR}^{i,(n+1)}).$
@@
$\hskip 41.25641pt\text{Compute }\mathbb{W}_{i}^{<2>}=\mathbb{Q}^{<2>}_{i}\mathbb{R}_{i}~\text{with }\mathbb{Q}^{<2>}_{i}\in\mathbb{R}^{r_{i-1}n_{i}\times r_{i}},~\mathbb{R}_{i}\in\mathbb{R}^{r_{i}\times r_{i}},$
@@
$\hskip 41.25641pt\mathbb{U}_{\rm LR}^{i,(n+1)}=(\mathbb{U}^{(n+1)}_{1},\ldots,\mathbb{U}^{(n+1)}_{i-1},\widetilde{\mathbb{U}}_{i}^{(n)},\mathbb{U}^{(n)}_{i+1},\ldots,\mathbb{U}^{(n)}_{d})$
$\hskip 24.18501pt\hookrightarrow\mathbb{U}_{\rm LR}^{i+1,(n+1)}=(\mathbb{U}^{(n+1)}_{1},\ldots,\mathbb{U}^{(n+1)}_{i-1},\mathbb{U}^{(n+1)}_{i},\widetilde{\mathbb{U}}^{(n)}_{i+1},\ldots,\mathbb{U}^{(n)}_{d})$,
$\hskip 41.25641pt\text{where }\mathbb{U}^{(n+1)}_{i}=\mathbb{Q}_{i},~\widetilde{\mathbb{U}}^{(n)}_{i+1}=\mathbb{R}_{i}\mathbb{U}^{(n)}_{i+1}$.
@@
$\hskip 28.45274pt\text{Compute }\mathbb{W}_{i,i+1}^{<2>}=\mathbb{P}_{i}^{<2>}\mathbb{\Sigma}_{i}\mathbb{Q}_{i}^{<1>}~\text{with }\mathbb{P}_{i}^{<2>}\in\mathbb{R}^{r_{i-1}n_{i}\times\widetilde{r}_{i}},~\mathbb{Q}_{i}^{<1>}\in\mathbb{R}^{\widetilde{r}_{i}\times n_{i+1}r_{i+1}},$